2. (a) Express 5 sinh x + cosh x in the form Ae^x+Be^{-x}, where A and B are integers.

(b) Solve the equation 5 sinh x + cosh x + 5 = 0, giving your answer in the form ln a, where a ∈ R.


1
Expert's answer
2021-05-25T17:27:50-0400

(a)  5 sinh x + cosh x

"=5.\frac{e^x-e^{-x}}{2}+\frac{e^x+e^{-x}}{2}"

"=\frac{5}{2}e^x-\frac{5}{2}e^{-x}+\frac{1}{2}e^x+\frac{1}{2}e^{-x} \\
=\frac{6}{2}e^x+\frac{-4}{2}e^{-x}"

"\boxed{A=3 , B=-2}"


(b) 5 sinh x + cosh x + 5 = 0

"=5.\frac{e^x-e^{-x}}{2}+\frac{e^x+e^{-x}}{2}+5=0"

let "e^x" = t

than

"\frac{5}{2}(t-\frac{1}{t})+\frac{1}{2}(t+\frac{1}{t})+5=0"

"\implies3t^2+5t-2=0"

"t=\frac{-5\pm\sqrt{49}}{6}"

"e^x=\frac{-5+7}{6}""\because""e^x>0" (always)

"\boxed{x=ln(\frac{1}{3})}"

here a="\frac{1}{3}"